# Optimal control of stochastic reaction networks with entropic control cost and emergence of mode-switching strategies

### v1

@@ II.3 Optimal control with KL cost
where ¯$\overline{\nabla}_{s_{r}}f(n):=f(n+s_{r})-f(n)$ is the discrete gradient (see Sec. A of the Supplementary Material [27] for the derivation). As the slope of $c_{KL}$ goes to infinity, $\frac{\partial}{\partial k_{r}}c_{KL}(k_{r},k_{r}^{0})\rightarrow-\infty$ for $k_{r}\rightarrow 0$, the nonnegativity constraint over $k$ is automatically satisfied. In fact, the exponential function in Eq. (13) ensures that the optimal coefficient $k^{\dagger}_{r}$ is non-negative.
Note that the KL control cost in Eq. (11) is weighted by the propensity function $h_{r}(n)$ for each reaction $r\in R$, which makes the control cost function depend on the current state $n$. It means that even if the reaction rate coefficient is the same, the control cost for a reaction $r$ is higher when the reaction occurs more frequently. A similar form of the control cost function has been used in quantifying the biological cost of cellular reactions [31].
@@
Although it is still an ordinary differential equation on a possibly infinite domain, linearity allows us to calculate the value function efficiently. Note that the second term on the right-hand side is the backward operator (generator) $\mathcal{L}_{k}$, which is the adjoint of the forward operator $\mathcal{L}^{\dagger}_{k}$ in the chemical master Eq. (II.1). Moreover, the special form of Eq. (15) allows us to have the following probabilistic representation for $Z_{t}(n,\beta)$:
@@
This representation enables us to compute the value function by evaluating the expectation of the utility function with respect to the uncontrolled reaction rate coefficient $k^{0}$. For simple reaction systems, we could obtain the analytical expression of the value function as demonstrated in the following sections. Even if it is not possible, efficient Monte Carlo sampling techniques can be used to estimate the expectation. It is worth noting that this representation of the value function can be evaluated in a time-forward manner, whereas the usual optimal control problem requires the time-backward calculation of the HJB equation due to the dynamic programming principle.
The linearization of the HJB equation and efficient computation of the optimal control problems for certain control cost functions was reported by Kappen [24] for diffusion processes and by Todorov [26] for discrete-time Markov chains. As Theodorou and Todorov [33] discussed, the linearization of the HJB equation is possible if the control cost function is given by KL divergence between path measures. In this case, the optimal control problem is related to an estimation (filtering) problem. Similar mathematical properties had been found in relation to the duality of control and inference [21, 22, 23]. Similar properties for Markov jump processes are identified in recent studies [34, 35, 36] as well as one of the earliest studies by Fleming [21]. In this paper, we develop the optimal control framework for stochastic RN inspired by these previous studies. In Sec. B of the Supplementary Material [27], we elaborate on the path measure perspective and why the KL cost function in Eq. (11) works.
II.4 First exit optimal control with KL cost
We have hitherto focused on the problem with a finite and fixed terminal time $T$. One can extend our approach to the problem with infinite time length $T\rightarrow\infty$. Since a stochastic reaction system often has absorbing states $\mathcal{N}_{\mathrm{abs}}\subset\mathcal{N}$, the stochastic process inevitably reaches one of these absorbing states after some time $T_{\mathrm{exit}}:=\min\{t\geq 0|n(t)\in\mathcal{N}_{\mathrm{abs}}\}$. For example, in the birth-death processes
the extinction state $n_{A}=0$ is absorbing. The extinction events are particularly important in biological problems because the extinction of some spieces in a chemical, organismal, or human population could alter the dynamics of the system qualitatively. Thus, controlling species into either survival or extinction has many applications, as mentioned in the Introduction. Moreover, biological control problems may not always have a prescribed end time $T$ because the goal of control is usually to achieve something rather than to do something until $T$. Thus, the first exist control is more essential than fixed-time control. Nevertheless, it has been less practiced because the first exist control is more involved technically.
We can consider the optimal control problem with such random terminal time $T_{\mathrm{exit}}$ and KL cost. The utility function is replaced with the sum of the terminal utility function $U_{\mathrm{exit}}(n)$ defined on the absorbing states $\mathcal{N}_{\mathrm{abs}}$, and the instantaneous utility function $U(n)$ defined on the non-absorbing states $\mathcal{N}_{\mathrm{non}}:=\mathcal{N}\backslash\mathcal{N}_{\mathrm{abs}}$, i.e.,
[…]
Then, the value function becomes time-independent $V(n,\beta)=V_{t}(n,\beta)$. If we assume the KL control cost, and the Cole–Hopf transformation $Z(n,\beta):=\exp(\beta V(n,\beta))$ yields
@@ II.4 First exit optimal control with KL cost
for $n\in\mathcal{N}_{\mathrm{non}}$, and $Z(n,\beta)=\exp\left(\beta U_{\mathrm{exit}}(n,\beta)\right)$ for $n\in\mathcal{N}_{\mathrm{abs}}$. Thus, the first exit optimal control problem is reduced to a linear algebraic problem. Similar results have been obtained for discrete-time Markov chains [26]. The probabilistic representations in Eqs. (16), (17) are also applicable.
@@
Via the Cole–Hopf transformation, the optimal solution could be cast into an eigenvalue problem (see Sec. C of the Supplementary Material [27]).
II.5 Controlling a random walker
Figure 2: The results of the minimum exit time problem for interacting random walkers on one-dimensional space. (a) The value function $V(x_{1},x_{2},\beta)$ plotted by color in the state space $(x_{1},x_{2})$ with different $\beta$. The dashed curves are its contours. The blue zigzag lines are the trajectories of $100$ independent simulations starting from $(10,11)$. (b) The value function $V(x_{1},x_{2},\beta)$ for fixed $x_{1}=0$ (the upper panel) and $x_{1}=10$ (the lower panel) with different $\beta$ indicated by the dots. The dashed lines are calculated with Eq. (25). The color code represents the value of $\beta$. (c) The gradient ¯$\overline{\nabla}_{2}V(x_{1},x_{2},\beta)=V(x_{1},x_{2}+1,\beta)-V(x_{1},x_{2},\beta)$ of the value function in $x_{2}$ direction for fixed $x_{1}=0$ and $x_{1}=10$. The dashed lines correspond to Eq. (25). The format of the panels is the same as in (b). The parameters are $M=31$, $N=2$, $k^{0}=0.1$.
@@ II.5 Controlling a random walker
$\underset{k(\cdot)}{\text{minimize}}\quad\beta\mathbb{E}_{k(\cdot)}\left[T_{\mathrm{exit}}\right]+\mathcal{D}_{KL}\left[\mathbb{P}_{k(\cdot)}\|\mathbb{P}_{k^{0}}\right],$ (22)
where the absorbing states are given by $\mathcal{N}_{\mathrm{abs}}=\{n\in\mathcal{N}|\hat{x}(n)=x^{*}\}$. Setting $U_{\mathrm{exit}}(n)=0$ and $U(n)=-1$, we have an equivalent maximization problem
$\underset{k(\cdot)}{\text{maximize}}\quad\beta\mathbb{E}_{k(\cdot)}\left[-T_{\mathrm{exit}}\right]-\mathcal{D}_{KL}\left[\mathbb{P}_{k(\cdot)}\|\mathbb{P}_{k^{0}}\right],$ (23)
Equation (17) shows that the value function $V(n,\beta)=V(\hat{x}(n),\beta)$ is equal to the cumulant-generating function of the exit time with parameter $-\beta$ and given by
@@
where $\Delta:=x^{*}-x$. Since the value function is linear in $|x^{*}-x|$, we have the piecewise constant optimal transition rates:
$k^{\dagger}_{(x,x+1)}(n,\beta)$ =$=\begin{cases}\kappa e^{\gamma(\beta)}&\text{if}~{}x<x^{*},\\ \kappa e^{-\gamma(\beta)}&\text{if}~{}x>x^{*},\end{cases}$ (28)
$k^{\dagger}_{(x,x-1)}(n,\beta)$ =$=\begin{cases}\kappa e^{-\gamma(\beta)}&\text{if}~{}x<x^{*},\\ \kappa e^{\gamma(\beta)}&\text{if}~{}x>x^{*}.\end{cases}$
When $x<x^{*}$, the transition rate $k^{\dagger}_{(x,x+1)}$ in the positive direction is higher than the uncontrolled one $\kappa$, while the transition rate $k^{\dagger}_{(x,x-1)}$ in the negative direction is lower than $\kappa$. We can also obtain the analytic solution for the standard control problem to maximize the average speed of the walker (see Sec. D of the Supplementary Material [27]).
II.6 Controlling interacting random walkers
@@
Among the abosrbing states above, we are interested in the single state $n^{*}\in\mathcal{N}_{\mathrm{abs}}$ such that the particles reach the goal $\hat{x}_{1}(n^{*})=\hat{x}_{2}(n^{*})=x^{*}$. We can formulate the problem as a first exit problem with the instantaneous utility $U(n)=-1$, and the terminal utility
@@
$=\mathbb{E}_{k^{0}}\left[\mathbb{E}_{k^{0}}\left[\exp(-\beta\max\{T^{1}_{\mathrm{exit}},T^{2}_{\mathrm{exit}}\})\big{|}\right.\right.$
$\quad\quad\left.\left.x_{1}(0)=x_{1},x_{1}(T^{1}_{\mathrm{exit}})=x^{*}\right]\right.$
$\quad\quad\left.x_{2}(0)=x_{2},x_{2}(T^{2}_{\mathrm{exit}})=x^{*}\right]$
@@
When the first particle has already reached the goal $x_{1}=0=x^{*}$, the problem is equivalent to the minimum time problem for the second particle only. Then, the value function satisfies $V(x^{*},x_{2},\beta)=-\gamma(\beta)|x^{*}-x_{2}|$, which is consistent with Eq. (33) (Fig. 2 (b), upper panel). Then, the gradient of $V(x^{*},x_{2},\beta)$ is constant (Fig. 2 (c), upper panel), which results in the constant optimal control.
@@
When the first particle is at $x_{1}=10$, the value $V(10,x_{2},\beta)$ as a function of $x_{2}$ has a gentle slope than $\gamma(\beta)$ to avoid collision with the first particle (Fig. 2 (b), lower panel). Especially when $\beta$ is small, collision avoidance is more important than early arrival. In this case, the value gradient can be positive, and the second particle moves away from the first particle, as seen in Fig. 2 (c) lower panel. On the other hand, when $\beta$ is large, the exit time becomes more important than the collision. Thus, the value gradient is always negative except at the collision point, i.e., at $x_{2}=10$.
II.7 Controlling survival in birth and death processes
@@
Figure 3: The results of the maximum exit time problem for Moran processes with $N=100$. (a) Dependence of the value function $V(n,\beta)$ on $\beta$ at $n_{A}=N/2$ with varying $\gamma$, the selective advantage of $A$ over $B$. $\gamma=1$ is the neutral case. Both the linear algebraic solution (colored points) and the analytical solution (colored curves) show divergence at $\beta=\beta_{c}$ (black dashed lines). (b) The relationship $\mathcal{E}(n,u)$ between the expected first exit time $u$ and the control cost starting from the state $n_{A}=N/2$. The format of the plot is the same as in (a). (c) The relationship $\mathcal{E}(n,u)/u$ between the expected first exit time and the control cost rate starting from the state $n_{A}=N/2$. The black dashed lines represent $\beta_{c}$ for each $\gamma$. The format of the plot is the same as in (a). (d, e) Dependence of the value function $V(n,\beta)$ (d) and its gradient ¯$\overline{\nabla}_{s_{1}}V(n,\beta)$ (e) on the initial state $n_{A}$ for different values of $\beta$ and $\gamma$. In each panel, the values of $\beta$ are sampled at equal intervals between 00 to $\beta_{c}$ and color-coded. (f) The optimally controlled stochastic trajectories from time 00 to $T=100$ for different $\beta$ and $\gamma$. In each panel, there are $100$ trajectories with initial conditions $n_{A}(0)=1$ for different values of $\beta$, which are color-coded. The parameters are $k_{1}^{0}=\gamma,k_{2}^{0}=1.0$.
@@
In the control of birth and death processes of cells or organisms without immigration, the extinction state $n=0$ is a natural absorbing state. In the context of biodiversity conservation, one has to avoid the extinction. Furthermore, a single species should not dominate the ecosystem. To address this population control problem, let us consider the following birth and death reactions involving two species $A$ and $B$ (Fig. 1 (b)):
@@
$\underset{k(\cdot)}{\text{maximize}}\quad\beta\mathbb{E}_{k(\cdot)}\left[T_{\mathrm{exit}}\right]-\mathcal{D}_{KL}\left[\mathbb{P}_{k(\cdot)}\|\mathbb{P}_{k^{0}}\right].$ (37)
The value function, i.e., the cumulant generating function of the extinction time, can be decomposed as
@@
where $K_{0}(n,\beta)$ and $K_{N}(n,\beta)$ are the cumulant-generating function of the first hitting time at 00 and $N$, respectively. They have an explicit expression [45] using the eigenvalues of submatrices of the transition rate matrix $M:=(-\omega_{n,n^{\prime}})_{n_{A},n_{A}^{\prime}\neq 0,N}\in\mathbb{R}^{(N-1)\times(N-1)}$ as follows:
$K_{0}(n,\beta)$ $:=\log\mathbb{E}_{k^{0}}\left[\exp(\beta T_{\mathrm{exit}})\delta_{n_{A}(T_{\mathrm{exit}}),0}|n(0)=n\right]$ (39)
$=\sum_{a=1}^{n_{A}}\log d^{0}(a)+\sum_{i=1}^{N-n_{A}-1}\log(x_{i}(n_{A})-\beta)$
@@
$K_{N}(n,\beta)$ $:=\log\mathbb{E}_{k^{0}}\left[\exp(\beta T_{\mathrm{exit,N}})\delta_{n_{A}(T_{\mathrm{exit}}),N}|n(0)=n\right]$ (40)
$=\sum_{a=n_{A}}^{N-1}\log b^{0}(a)+\sum_{j=1}^{n_{A}-1}\log(y_{j}(n_{A})-\beta)$
@@
As a function of $n_{A}$, the value function $V(n,\beta)$ for each $\beta$ and $\gamma$ has a single peak, which we designate by $n_{A}^{*}$ (Fig. 3 (d)). The optimal control steers the system to stay around $n_{A}^{*}$. As $\beta$ increases, the value gradient becomes steep around $n_{A}^{*}$ (Fig. 3 (e)) so that the controlled trajectories do not leave there even after a long time (Fig. 3 (f)).
@@
When $\beta$ is close to $\beta_{c}$, the width of the peak around $n_{A}^{*}$ decreases as $\gamma$ increases, and there emerges a region with a shallow gradient ¯$\overline{\nabla}_{s_{1}}V(n,\beta)\approx 0$. In this region, the optimal control $k^{\dagger}_{r}(n)=k^{0}_{r}\exp(\overline{\nabla}_{s_{r}}V(n,\beta))$ become close to the uncontrolled rate $k^{0}_{r}$. The emergence of the zero-gradient region indicates that the optimal control strategy switches between ON and OFF modes depending on the current state.
@@
The OFF mode region appears if $\gamma$ is high and $N$ is moderately large (see the Fig. S1 in Sec. E of the Supplementary Material [27]), implying that the OFF mode is attributed to the difficulty of controlling an exponentially growing population. If both $\gamma$ and $N$ are large, the uncontrolled system has a stronger and less stochastic driving force towards $n_{A}=N$ as $n_{A}$ increases. Thus, the optimal control keeps the population away from $n_{A}=N$ by forcing it close to 00, which is the ON mode around the peak $n_{A}^{*}$. Once a large fluctuation drives the population to leave the peak, the additional cost to bring it back to the peak does not pay for its success rate, leading to the OFF mode, i.e., do nothing, in the intermediate region. Finally, the control turns ON again near $n_{A}=N$ to hang on there. Such a spontaneous emergence of hierarchical control may not be identified within the LQG approximation, highlighting the importance of taking into account the unique properties of RN.
II.8 Controlling epidemic outbreak

### v2

@@ II.5 Optimal control with KL cost
(see Appendix A for the derivation). As the slope of $C_{KL}$ goes to infinity, $\frac{\partial}{\partial k_{r}}C_{KL}(n,k)\rightarrow$-$\infty$ for $k_{r}\rightarrow 0$, the nonnegativity constraint over $k$ is automatically satisfied. In fact, the exponential function in Eq. (32) ensures that the optimal coefficient $k^{\dagger}_{r}$ is non-negative.
@@
Although it is still an ordinary differential equation on a possibly infinite domain, the linearity allows us to calculate the value function efficiently. Note that the second term on the right-hand side is the Kolmogorov backward operator $\mathcal{L}_{k}$, which is the adjoint of the Kolmogorov forward operator $\mathcal{L}^{\dagger}_{k}$ in the chemical master equation (Eq. (II.1)). Moreover, the special form of Eq. (34) allows us to have the following probabilistic representation for $Z_{t}(n,\beta)$:
@@
This representation enables us to compute the value function by evaluating the expectation of the utility function with respect to the uncontrolled reaction rate coefficient $k^{0}$. As demonstrated in the following sections, for simple reaction systems, we can obtain the analytical expression of the value function. Even if an analytical expression is inaccessible, efficient Monte Carlo sampling techniques can be used to estimate the expectation. It is worth noting that this representation of the value function can be evaluated in a time-forward manner, whereas the usual optimal control problem requires the time-backward calculation of the HJB equation due to the dynamic programming principle.
The linearization of the HJB equation and efficient computation of the optimal control problems for certain control cost functions was reported by Kappen [28] for diffusion processes and by Todorov [30] for discrete-time Markov chains. As Theodorou and Todorov [40] discussed, the linearization of the HJB equation is possible if the control cost function is given by KL divergence between path measures. In this case, the optimal control problem is related to an estimation (filtering) problem. Similar mathematical properties had been found in relation to the duality of control and inference [25, 26, 27]. Similar properties for Markov jump processes are identified in recent studies [41, 42, 43] as well as one of the earliest studies by Fleming [25]. In this paper, we develop the optimal control framework for stochastic RN inspired by these previous studies. In Appendix C, we elaborate on the path measure perspective and why the KL cost function in Eq. (31) works. See Appendix E for further comparison with existing literature.
The same trick applies to the first exit problems. The Cole–Hopf transformation $Z(n,\beta):=\exp(\beta V(n,\beta))$ yields
@@
for $n\in\mathcal{N}_{\mathrm{non}}$, and $Z(n,\beta)=\exp\left(\beta U_{\mathrm{exit}}(n,\beta)\right)$ for $n\in\mathcal{N}_{\mathrm{abs}}$. The equations can be written in the matrix-vector form:
$\bm{0}=\left(\beta\mathop{\rm diag}\nolimits\bm{U}+\Omega^{0}_{\mathrm{non},\mathrm{non}}\right)\bm{Z}_{\mathrm{non}}+\Omega^{0}_{\mathrm{non},\mathrm{abs}}\bm{Z}_{\mathrm{abs}},$ (38)
where $\bm{U}=\{U(n)\}$ is a $|\mathcal{N}_{\mathrm{non}}|$-dimensional vector and the matrices $\Omega^{0}_{\mathrm{non},\mathrm{non}}\in\mathbb{R}^{|\mathcal{N}_{\mathrm{non}}|\times|\mathcal{N}_{\mathrm{non}}|}$ and $\Omega^{0}_{\mathrm{non},\mathrm{abs}}\in\mathbb{R}^{|\mathcal{N}_{\mathrm{non}}|\times|\mathcal{N}_{\mathrm{abs}}|}$ are given by the the uncontrolled transition rate $\omega^{0}_{n,n^{\prime}}:=\sum_{r\in R}k^{0}_{r}h_{r}(n)\delta_{n+s_{r},n^{\prime}}$ as in Eq. (4). From the boundary condition, $\bm{Z}_{\mathrm{abs}}$ is known. We can solve the above equation for $\bm{Z}_{\mathrm{non}}$. Thus, the first exit optimal control problem is reduced to a linear algebraic problem. Similar results have been obtained for discrete-time Markov chains [30]. The probabilistic representations in Eqs. (35), (36) are also applicable.
@@
Via the Cole–Hopf transformation, the optimal solution could be cast into an eigenvalue problem (see Appendix D).
II.6 Controlling a random walker
Figure 3: The results of the minimum exit time problem for interacting random walkers on one-dimensional space. (a) The value function $V(x_{1},x_{2},\beta)$ plotted by color in the state space $(x_{1},x_{2})$ with different $\beta$. The dashed curves are its contours. The blue zigzag lines are the trajectories of $100$ independent simulations starting from $(10,11)$. (b) The value function $V(x_{1},x_{2},\beta)$ for fixed $x_{1}=0$ (the upper panel) and $x_{1}=10$ (the lower panel) with different $\beta$ indicated by the dots. The dashed lines are calculated with Eq. (43). The color code represents the value of $\beta$. (c) The directional derivative $\overline{\nabla}_{2}V(x_{1},x_{2},\beta)=V(x_{1},x_{2}+1,\beta)-V(x_{1},x_{2},\beta)$ of the value function in $x_{2}$ direction for fixed $x_{1}=0$ and $x_{1}=10$. The dashed lines correspond to Eq. (43). The format of the panels is the same as in (b). The parameters are $M=31$, $N=2$, $k^{0}=0.1$.
@@ II.6 Controlling a random walker
$\underset{k(\cdot)}{\text{minimize}}\quad\mathbb{E}_{k(\cdot)}\left[\beta T_{\mathrm{exit}}+\int_{0}^{T_{\mathrm{exit}}}C(n(t),k(t))dt\right],$ (40)
where the absorbing states are given by $\mathcal{N}_{\mathrm{abs}}=\{n\in\mathcal{N}|\hat{x}(n)=x^{*}\}$. By setting $U_{\mathrm{exit}}(n)=0$ and U(n)=1U(n)=-1 in Eq. (16), we have the following maximization problem which is equivalent to Eq. (40) having the same form as Eq. (7):
$\underset{k(\cdot)}{\text{maximize}}\quad\mathbb{E}_{k(\cdot)}\left[-\beta T_{\mathrm{exit}}-\mathcal{C}(n(\cdot),k(\cdot))\right].$ (41)
If we adopt the KL control cost function, Eq. (36) shows that the value function $V(n,\beta)=V(\hat{x}(n),\beta)$ is equal to the cumulant-generating function of the exit time with parameter $\beta$ and given by
@@
where $\Delta:=x^{*}-x$. Since the value function is linear in $|x^{*}-x|$, we have the piecewise constant optimal transition rates by using Eqs. (32) and (43):
$k^{\dagger}_{(x,x+1)}(n,\beta)$ $κeγ(β)ifx>x,\displaystyle$= (46)
[…]
When $x<x^{*}$, the transition rate $k^{\dagger}_{(x,x+1)}$ in the positive direction is higher than the uncontrolled one $\kappa$, while the transition rate $k^{\dagger}_{(x,x-1)}$ in the negative direction is lower than $\kappa$. We can also obtain the analytic solution for the standard control problem to maximize the average speed of the walker (see Appendix F).
II.7 Controlling interacting random walkers
@@
Among the abosrbing states above, we are interested in the single state $n^{*}\in\mathcal{N}_{\mathrm{abs}}$ such that the particles reach the goal $\hat{x}_{1}(n^{*})=\hat{x}_{2}(n^{*})=x^{*}$. We can formulate the problem as a first exit problem with the instantaneous utility U(n)=1U(n)=-1, and the terminal utility
@@
[…]
$\quad\quad\big|\left.\left.x_{1}(0)=x_{1},x_{1}(T^{1}_{\mathrm{exit}})=x^{*}\right]\right$.
$\quad\quad\big|\left.x_{2}(0)=x_{2},x_{2}(T^{2}_{\mathrm{exit}})=x^{*}\right]$
@@
When the first particle has already reached the goal $x_{1}=0=x^{*}$, the problem is equivalent to the minimum time problem for the second particle only. Then, the value function satisfies $V(x^{*},x_{2},\beta)=-\gamma(\beta)|x^{*}-x_{2}|$, which is consistent with Eq. (51) (Fig. 3 (b), upper panel). Then, the derivative of $V(x^{*},x_{2},\beta)$ is constant (Fig. 3 (c), upper panel), which results in the constant optimal control.
@@
When the first particle is at $x_{1}=10$, the value $V(10,x_{2},\beta)$ as a function of $x_{2}$ has a gentle slope than $\gamma(\beta)$ to avoid collision with the first particle (Fig. 3 (b), lower panel). Especially when $\beta$ is small, collision avoidance is more important than early arrival. In this case, the derivative of the value can be positive, and the second particle moves away from the first particle, as seen in Fig. 3 (c) lower panel. On the other hand, when $\beta$ is large, the exit time becomes more important than the collision. Thus, the derivative is always negative except at the collision point, i.e., at $x_{2}=10$.
II.8 Controlling survival in birth and death processes
@@
Figure 4: The results of the maximum exit time problem for Moran processes with $N=100$. (a) Dependence of the value function $V(n,\beta)$ on $\beta$ at $n_{A}=N/2$ with varying $\gamma$, the selective advantage of AA over $B$. $\gamma=1$ is the neutral case. Both the linear algebraic solution (colored points) and the analytical solution (colored curves) show divergence at $\beta=\beta_{c}$ (black dashed lines). (b) The relationship $\mathcal{E}(n,u)$ between the expected first exit time $u$ and the control cost starting from the state $n_{A}=N/2$. The format of the plot is the same as in (a). (c) The relationship $\mathcal{E}(n,u)/u$ between the expected first exit time and the control cost rate starting from the state $n_{A}=N/2$. The black dashed lines represent $\beta_{c}$ for each $\gamma$. The format of the plot is the same as in (a). (d, e) Dependence of the value function $V(n,\beta)$ (d) and its derivative $\overline{\nabla}_{s_{1}}V(n,\beta)$ (e) on the initial state $n_{A}$ for different values of $\beta$ and $\gamma$. In each panel, the values of $\beta$ are sampled at equal intervals between 0 to $\beta_{c}$ and color-coded. (f) The optimally controlled stochastic trajectories from time 0 to $T=100$ for different $\beta$ and $\gamma$. In each panel, there are 100100 trajectories with initial conditions $n_{A}(0)=1$ for different values of $\beta$, which are color-coded. The parameters are $k_{1}^{0}=\gamma,k_{2}^{0}=1.0$.
Figure 5: Comparison of maximum exit time solution for Moran processes with different control cost functions. (a–c) The value functions with the quadratic control cost function $C_{quad}$ are shown as in Fig. 4 (a-c). (d–f) The value functions and optimally controlled trajectories with the quadratic $C_{quad}$, weighted KL $C_{wKL}$ ($w_{1}=\infty,w_{2}=1$), KL $C_{KL}$, and weighted KL $C_{wKL}$ ($w_{1}=1,w_{2}=\infty$) control cost functions. The parameters are $N=100,k_{1}^{0}=1.5,k_{2}^{0}=1.0$.
@@
In the control of birth and death processes of cells or organisms without immigration, the extinction state $n=0$ is a natural absorbing state. In the context of biodiversity conservation, one has to avoid extinction. Furthermore, a single species should not dominate the ecosystem. To address this population control problem, let us consider the following birth and death reactions involving two species $A$ and $B$ (Fig. 1 (b)):
@@
$\underset{k(\cdot)}{\text{maximize}}\quad\mathbb{E}_{k(\cdot)}\left[\beta T_{\mathrm{exit}}-\mathcal{C}(n(\cdot),k(\cdot))\right].$ (55)
Assuming the KL control cost, we can derive an analytical solution. Using the fact that the terminal state $n_{A}(T_{\mathrm{exit}})$ is either 0 or $N$, the value function, i.e., the cumulant generating function of the extinction time, can be decomposed as
@@
$=\log\mathbb{E}_{k^{0}}\left[\exp(\beta T_{\mathrm{exit}})\right$.
$\quad\quad\quad\left.(\delta_{n_{A}(T_{\mathrm{exit}}),0}+\delta_{n_{A}(T_{\mathrm{exit}}),N})|n(0)=n\right]$
@@
where
$K_{0}(n,\beta)$ $=\log\mathbb{E}_{k^{0}}\left[\exp(\beta T_{\mathrm{exit}})\delta_{n_{A}(T_{\mathrm{exit}}),0}|n(0)=n\right],$ (57)
$K_{N}(n,\beta)$ $=\log\mathbb{E}_{k^{0}}\left[\exp(\beta T_{\mathrm{exit}})\delta_{n_{A}(T_{\mathrm{exit}}),N}|n(0)=n\right]$
are the cumulant-generating functions of the first hitting time at 0 and $N$, respectively. They have an explicit expression [52] using the eigenvalues of submatrices of the transition rate matrix $M:=(-\omega_{n,n^{\prime}})_{n_{A},n_{A}^{\prime}\neq 0,N}\in\mathbb{R}^{(N-1)\times(N-1)}$ as follows:
$K_{0}(n,\beta)$ $=\sum_{a=1}^{n_{A}}\log d^{0}(a)+\sum_{i=1}^{N-n_{A}-1}\log(x_{i}(n_{A})-\beta)$ (58)
@@
$K_{N}(n,\beta)$ $=\sum_{a=n_{A}}^{N-1}\log b^{0}(a)+\sum_{j=1}^{n_{A}-1}\log(y_{j}(n_{A})-\beta)$ (59)
@@
As a function of $n_{A}$, the value function $V(n,\beta)$ for each $\beta$ and $\gamma$ has a single peak, which we designate by $n_{A}^{*}$ (Fig. 4 (d)). The optimal control steers the system to stay around $n_{A}^{*}$. As $\beta$ increases, the derivative of the value function becomes steep around $n_{A}^{*}$ (Fig. 4 (e)) so that the controlled trajectories do not leave there even after a long time (Fig. 4 (f)).
@@
When $\beta$ is close to $\beta_{c}$, the width of the peak around $n_{A}^{*}$ decreases as $\gamma$ increases, and there emerges a region with a shallow slope $\overline{\nabla}_{s_{1}}V(n,\beta)\approx 0$. In this region, the optimal control $k^{\dagger}_{r}(n)=k^{0}_{r}\exp(\overline{\nabla}_{s_{r}}V(n,\beta))$ become close to the uncontrolled rate $k^{0}_{r}$. The emergence of the shallow slope region indicates that the optimal control strategy switches between ON and OFF modes depending on the current state.
@@
The OFF mode region appears if $\gamma$ is high and $N$ is moderately large (see Appendix G), implying that the OFF mode is attributed to the difficulty of controlling an exponentially growing population. If both $\gamma$ and $N$ are large, the uncontrolled system has a stronger and less stochastic driving force towards $n_{A}=N$ as $n_{A}$ increases. Thus, the optimal control keeps the population away from $n_{A}=N$ by forcing it close to 0, which is the ON mode around the peak $n_{A}^{*}$. Once a large fluctuation drives the population to leave the peak, the additional cost of bringing it back to the peak does not compensate for its success rate, leading to the OFF mode, i.e., doing nothing, in the intermediate region. Finally, the control turns ON again near $n_{A}=N$ to hang on there. Such a spontaneous emergence of hierarchical control may not be identifiable within the LQG approximation, highlighting the importance of considering the unique properties of RN.
The result explained so far used the KL control cost problem. The existence of a finite cost rate to prevent ultimate extinction and the emergence of the mode-switching control strategy could be attributed to either the special choice of the control cost function, the problem itself (maximizing extinction time for Moran processes), or both. To reveal the effect of the control cost function, we numerically solved the optimal control problem with two types of non-KL control cost functions. The first control cost function is the quadratic control cost $C_{quad}$ given by Eq. (21). While the quadratic and KL control cost functions are the same up to the second order in $k-k^{0}$ around $k=k^{0}$, they behave differently in regions far from $k^{0}$. The second class of control cost functions is the weighted KL control cost given by
$C_{wKL}(n,k)=\sum_{r\in R}w_{r}h_{r}(n)\left(k_{r}\log\frac{k_{r}}{k_{r}^{0}}-k_{r}+k_{r}^{0}\right),$ (60)
where $w_{r}>0$ is the weighting factor for each reaction. When the weighting factor for reaction $r$ is infinite, i.e., $w_{r}=+\infty$, the reaction rate coefficient $k_{r}$ is constrained to be $k_{r}^{0}$.
We obtained the optimal solutions by numerically solving the nonlinear HJB Eq. (28). As shown in Fig. 5 (a-c), the value function $V(n,\beta)$ as a function of $\beta$ and the cost-exit time tradeoff curves $\mathcal{E}(n,u)$ for the quadratic control cost are qualitatively similar to that of KL control cost. The existence of a critical cost rate, $\beta_{c}$, that prevents ultimate extinction can be confirmed. Furthermore, as shown in Fig. 5 (d-f), the value functions as a function of $n_{A}$ with the various non-KL control costs are qualitatively similar to those of the KL control cost. Despite the quantitative difference, the derivative of the value function always exhibits the OFF mode region. Consequently, the simulated controlled trajectories show the mode-switching behavior. Therefore, the emergence of the mode-switching strategy is not due to the nature of the KL control cost, but rather an intrinsic feature of the control problem for Moran processes.
II.9 Controlling epidemic outbreak